How many moles of aluminum oxide (Al2O3) are produced when 4.0 moles of aluminum (Al) are combined with Oxygen gas (O2)?
1
Expert's answer
2020-11-25T02:24:24-0500

4Al+3O2=2Al2O3

n (Al2O3) = 1/2 x n (Al) = 1/2 x 4 =2 moles